An order is written for 15 grams of a 5% lidocaine topical cream. Only 3% concentration and 6% concentration are available. How much of the 3% concentration must be added to the 6% concentration to fill this order?
Answer by Alan3354(69443)

An order is written for 15 grams of a 5% lidocaine topical cream. Only 3% concentration and 6% concentration are available. How much of the 3% concentration must be added to the 6% concentration to fill this order?
------------
t = amount of 3%
s = amount of 6%
---
t + s = 15
3t + 6s = 15*5